Consider 6 resistors 10,12,18,20,24 and 32 connected in parallel to a 64 V battery
1
Expert's answer
2020-03-09T11:15:57-0400

In parallel voltage are same across each resistance so

Current can be calculated as

I1 = 64/10 = 6.4 A

I2 = 64/12 = 5.34 A

I3 = 64/18 = 3.56 A

I4 = 64/20 = 3.2 A

I5 = 64/24 = 2.67 A

I6 = 64/32 = 2A